Displacement of a partial is given by S=3t^3+7t^2+5t+8.acceleration of the partical is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Sanchita
Displacement (s) = 3 t³ + 7t² +5t+8
Velocity (v) = ds/dt = 9t² + 14t +5
Acceleration (a) = dv/dt = d²s/dt² = 18t +14
Regards
Arun (askIITians forum expert)